3) For the given probability distribution, find the mean and standard deviation
Answer #1

Solution :

Mean = \mu = \sum X * P(X)

= 1 *0.12 + 2 * 0.23 + 3 * 0.35 + 4 * 0.18+5*0.12

= (0.12 +0.46 +1.05 + 0.72+0.6)

mean= 2.95

Standard deviation = \sigma

=\sqrt\sumX 2 * P(X) - \mu 2

=  \sqrt[12 *0.12 + 22 * 0.23 + 32 * 0.35 + 42 * 0.18+52*0.12] -2.95 2

= \sqrt [(0.12 +0.92 +3.15 +2.88+3 ) )] -8.7025

=\sqrt 10.07 -8.7025

=\sqrt1.3675

\sigma=1.1694